3. Z-test for the population mean u (o known) Hosmer Winery (on Cayuga Lake, New York) manufactures and bottles wine for sale• Calculate the p-value. TABLE A Standard Normal Cumulative Probabilities (continued) .00 02 03 .04 .05 06 .07 .01 .08 .09 2.
Answer #1

(a) The Hypothesis

H0: \mu = 753

Ha: \mu \neq 753

_________________

(b) By the property of Normality, since the population from which the sample comes from is given to be normal, therefore the distribution of the sample means can also be considered to be normal, regardless of the size of N. This is called the First Known Property.

___________________

(c) \bar{x} = (745 + 744 + 750 + 745 + 756 + 753 + 748 + 751) / 8 = 5992/8 = 749

The Test Statistic:

Z = \frac{\bar{x}-\mu }{\frac{\sigma}{\sqrt{n}}} = \frac{749-753}{\frac{4.3}{\sqrt{8}}} = -2.63

The p value (2 tail) : Since z is -ve, we find the p value for -2.63 and then multiply the value by 2 = 0.0043 * 2 = 0.0086

Since p value is < 0.05, reject H0.

The Conclusion: There is sufficient evidence at the 95% significance level to conclude that the amount of wine the machine dispenses is different from 253 ml.

__________________________________

(d) The Zcritical (2 tail) for \alpha = 0.05, is 1.96

The Confidence Interval is given by \bar{x} \pm ME, where

ME = Zcritical * \frac{\sigma}{\sqrt{n}} = 1.96 * \frac{4.3}{\sqrt{8}} = 2.98

The Lower Limit = 749 - 2.98 = 746.02

The Upper Limit = 749 + 2.98 = 751.98

The Confidence Interval is (746.02 , 751.98)

The confidence interval does not contain the value of 753, and hence we can reject the null hypothesis that mean is equal to 753 ml, and conclude that the mean is different from 753 ml.

________________________________